[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat May 20, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat May 20 18:31:42 UTC 2017


HSU Shift Log for Sat May 20 2017 14:31:35 GMT-0400 (EDT)

Shift: H

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: GOES-16 ingest working fine.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170520/5caa3541/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list